OpenStudy (anonymous):

If I sent five 2 oz. letters and three 3 oz. letters by first class mail that cost me $5.80. Sending three 2 oz. letters and five 3 oz. letters cost me $6.20 what would the cost be for one 2 oz. and one 3 oz. letter?

OpenStudy (johnweldon1993):

System of equations Let 2 oz letters be represented as A And let 3oz letters be represented as B \[\large 5A + 3B = 5.80\] \[\large 3A + 5B = 6.20\] Once we solve for A and B in that system, we can use them to find out \[\large A + B = ?\]

OpenStudy (johnweldon1993):

Good here? :)

OpenStudy (anonymous):

thank you so much I got .65 for one and .85 for the other I appreciate your help